OpenStudy (anonymous):

y=(3x-7)/(2x+4). What is dy/dx?

OpenStudy (owlcoffee):

Remember the derivative of a product: \[(f.g)'(x) = f'(x)g(x)+f(x)g'(x)\]

OpenStudy (anonymous):

(f/g)'(x) = [f'(x)g(x)-f(x)g'(x)]/[g(x)^2] The quotient rule. I have worked it through a bunch of times but I am not getting the right answer

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=3x-7,f'(x)=3,g(x)=2x+4,g'(x)=2\]

OpenStudy (owlcoffee):

oh... It was a quotient... I'll help you. So we have: \[y= \frac{ (3x-7) }{ (2x+4) }\] so then y' must be: \[y'=\frac{ (3)(2x+4)-(3x-7)(2) }{ (2x+4)^{2} }\] All you have to do from here on is simplify that equation.

OpenStudy (anonymous):

I get 26/[(2x+4)^2] but webassign is telling me that its wrong

OpenStudy (owlcoffee):

True, see we have (2x+4) ^2 on the denominator, and a (2x+4) multiplying in the numerator, so we can simplify them, ending up with: \[y'= \frac{ 3-(3x-7)(2) }{ (2x+4) }\]

OpenStudy (anonymous):

I don't think you can do that

OpenStudy (anonymous):

answer is 13/[2(x+2)^2] but I still don't know how to get this answer

OpenStudy (anonymous):